At the gym club there is 2 boys for every 3 girls there is 50 children how many girls are there?

Answer by checkley79(3341)

2M+3M=50
5M=50
M=50/5
M=10 THIS IS THE MULTIPLIER FOR THE RATIOS.
2*10=20 BOYS.
3*10=30 GIRLS
PROOF"
20+30=50
50=50